OpenStudy (anonymous):

a. Find the ratio of the corresponding altitudes of two similar triangles in which the ratio of the corresponding sides is 2:3. b. Find the ratio of the areas of two similar triangles in which the ratio of the corresponding sides 2:3.

OpenStudy (amistre64):

is there pictures or are we sposed to use our imagination?

OpenStudy (anonymous):

There are no pictures... I guess we supposed to use our imagination.

OpenStudy (amistre64):

is that 2:3 a ration of triangle to triangle? or is it one side of a triangle to another side of the same triangle?

OpenStudy (amistre64):

are these right tris or any tris will work?

OpenStudy (anonymous):

see area is always proportional to the square of the ratio. so area will be at a ratio of 4:9

OpenStudy (anonymous):

I think any triangle will work, I'm not sure I copied the question exactly the way it appears on my sheet. I think its talking about sides that correspond.

OpenStudy (anonymous):

Dipankarstudy I'm confused at what you just said to be honest, I kinda bad at math, I missed this whole topic

OpenStudy (amistre64):

2:3 --- 4:6 so if we have a tri with sides 2 and 3; we can make another with sides 4 and 6 right?

OpenStudy (amistre64):

gonna have to take a min and brush up on that terminology :)

OpenStudy (amistre64):

ok..... the sides have a 2/3 ratio....got it; let me draw a pic

OpenStudy (anonymous):

see for a similar triangle the sides are in proportion. see the attach ment... where the triangles are similar area of a triangle=base*hight/2; area abc 0.5*bc*ad 2*2 ------- = -------- = ----- =4/9 area def .5*fg*eh 3*3 ad and eh are also will be in the same ratio as the trin 'abd' and 'efh' are similar....

OpenStudy (amistre64):

this would be similar tris with a 2:3 ratio

OpenStudy (anonymous):

i think u have understood.....

OpenStudy (anonymous):

To be honest Im very confused the question seems to be a short answer with the space provided, and not that complicated.

OpenStudy (amistre64):

gotta work up to the answer; you wanna know the stuff right?

OpenStudy (anonymous):

Yea, but Im very confused and don't understand the question at all. The ratio part and the corresponding sides is really confusing me

OpenStudy (amistre64):

we can determine the ration of altitudes by comparing the side*sin ratio

OpenStudy (amistre64):

the ration just means that corresponding sides should equal 2/3 when written one on top of the other..

OpenStudy (anonymous):

Oh okay

OpenStudy (amistre64):

i dont know why i keep sticking that n onto my ratio wird lol

OpenStudy (amistre64):

is the 2sin(30) the 3sin(30) a ration of 2/3 ?

OpenStudy (amistre64):

in other words: is: 2 sin(30) ------- = 2/3 ? 3 sin(30)

OpenStudy (anonymous):

No it isn't I think -_-